OpenStudy (anonymous):

Find the distance between these points. A (5, 8), B(-3, 4) AB =

OpenStudy (anonymous):

Square root of \[(-3-5)^2 + (4-8)^2 = \sqrt{64 + 16} = \sqrt{80}\]

OpenStudy (anonymous):

on a side note, your instructor might want the answer in simplest radical form: \[\huge \sqrt{80}=\sqrt{(16)(5)}=\sqrt{16}\sqrt{5}=4\sqrt{5} \]

OpenStudy (anonymous):

Yes, we can solve it further.. Thanks for solving further...

OpenStudy (anonymous):

please allow me to add...attachment

OpenStudy (anonymous):

You are saying to whom??

OpenStudy (anonymous):

lorataylor please allow me to attach an explanation

OpenStudy (anonymous):

please bear with me my scanner is going very slow...

OpenStudy (anonymous):

after a whole century it uploaded anyhow the side note got cut off, it says: sqrt( )=( )^ 1/2 so sqrt(c^2)=(c^2)^1/2 =c^2*(1/2) =c^ 2/2 =c^1 =c

OpenStudy (anonymous):

What is it mashe??/

OpenStudy (anonymous):

I'm sorry I do not understand what do you mean by your question waterineyes? i uploaded the pdf "distance.pdf" you may click and view it.

OpenStudy (anonymous):

No no not the attachment. but the after a whole century it uploaded.........What is it??

OpenStudy (anonymous):

I said it took a whole century b/c my scanner took forever to upload the distance.pdf that i wanted lorataylor to take a look at but it seems hardly anyone cares about math the way I do and i wish people could really understand the question they are asking (sometimes people can't solve it b/c they don't know what they are being asked)...so lorataylor your question could have been phrased "Find the length of the hypotenuse of the right triangle with coordinates A(5,8), B(-3,4) and C=(5,4)"